# Traveling supersolid stripe patterns in spin-orbit-coupled Bose-Einstein condensates

### v1

@@ III.1 Laboratory frame
where $\mu$ is the chemical potential in the laboratory frame, and $\Psi_{0}$ is a function satisfying periodic boundary conditions in all three spatial directions. Substituting the traveling-wave Ansatz (7) into the Gross-Pitaevskii equation (1) and transforming to the comoving frame coordinates $\bm{r}^{\prime}=\bm{r}-\bm{v}t$ and $t^{\prime}=t$ (note that throughout this work, primed quantities refer to the comoving frame), we obtain the following time-independent equation for $\Psi_{0}$:
$\begin{split}&{}h_{\mathrm{SO}}\Psi_{0}+g_{dd}\left(\Psi_{0}^{\dagger}\Psi_{0}\right)\Psi_{0}+g_{ss}\left(\Psi_{0}^{\dagger}\sigma_{z}\Psi_{0}\right)\sigma_{z}\Psi_{0}\\ &{}=(\mu+\bm{v}\cdot\bm{p})\Psi_{0}\,,\end{split}$ (8)
@@
where now $\bm{p}=-\mathrm{i}\hbar\nabla_{\bm{r}^{\prime}}$. For $\bm{v}=0$, Eq. (8) reduces to the standard time-independent Gross-Pitaevskii equation, whose solutions represent the stationary states of the condensate, including the ground state. The $\bm{v}$-dependent term on the right hand side was previously included in Ref. [61] to construct supercurrent states, though that study considered only uniform plane-wave configurations. Note that Eq. (8) corresponds to the stationarity condition of the functional
$E_{\mathrm{GC}}^{\prime}=E-\mu N-\bm{v}\cdot\langle\bm{p}\rangle\,,$ (9)
which is the grand canonical energy in the comoving frame (as detailed below). Therefore, $\mu$ and $\bm{v}$ can be interpreted as Lagrange multipliers, determined by fixing the values of the particle number $N$ and the momentum $\langle\bm{p}\rangle$.
III.2 Comoving frame
$h_{\mathrm{SO}}^{\prime}=h_{\mathrm{SO}}-v_{x}\hbar k_{R}\sigma_{z}\,.$ (11)
This extra term can be understood by noting that in the comoving frame the Raman lasers move with velocity $-\bm{v}$. As a result, the frequency of the light field experienced by the atoms undergoes a Doppler shift of $-2v_{x}k_{R}$, which leads to the corresponding shift in the Raman detuning $\delta_{R}$ by the same amount [60, 78].
In the comoving frame, the periodic wave pattern is stationary, while the superfluid background is boosted with velocity $-\bm{v}$. This configuration can be described by a stationary wave function of the form
$\Psi^{\prime}(\bm{r}^{\prime},t^{\prime})=\mathrm{e}^{-\mathrm{i}\mu^{\prime}t^{\prime}/\hbar}\mathrm{e}^{-\mathrm{i}m\bm{v}\cdot\bm{r}^{\prime}/\hbar}\Psi_{0}(\bm{r}^{\prime})\,.$ (12)
Substituting this Ansatz into Eq. (10) and setting $\mu^{\prime}=\mu+mv^{2}/2$, one can easily recover Eq. (8). This confirms that the function $\Psi_{0}$ introduced in Eq. (12) is identical to the one used in the traveling-wave Ansatz (7). Additionally, the relation between the energies in the two frames is given by
$E^{\prime}=E-\bm{v}\cdot\langle\bm{p}\rangle+\frac{Nmv^{2}}{2}\,.$ (13)
Taking the derivative of both sides of this equation with respect to $N$, we see that $\mu^{\prime}$ is indeed the chemical potential in the comoving frame. Furthermore, the grand canonical energy $E_{\mathrm{GC}}^{\prime}=E^{\prime}-\mu^{\prime}N$ reproduces Eq. (9) above.
III.3 Traveling stripe Ansatz
The full condensate order parameter in the supersolid stripe phase can be expressed as an expansion of the form [76, 58]
$\Psi_{0}(\bm{r}^{\prime})=\sqrt{\bar{n}}\sum_{\text{{\hbox{\bar{m}}} odd}}\tilde{\Psi}_{\bar{m}}\mathrm{e}^{\mathrm{i}\bar{m}k_{1}x^{\prime}}\,.$ (14)
@@ III.3 Traveling stripe Ansatz
In this expression, the $\tilde{\Psi}_{\bar{m}}$’s are two-component expansion coefficients, and the sum runs over both positive and negative odd integers. The presence of the $\bar{m}=\pm 1$ terms in the Ansatz indicates the simultaneous occupation of the two minima of the single-particle spectrum. In contrast, the $|\bar{m}|>1$ harmonics arise due to the nonlinearity of the Gross-Pitaevskii equation (8) that $\Psi_{0}$ satisfies. Note that the order parameter (14) is $\pi/k_{1}$-antiperiodic (and therefore $2\pi/k_{1}$-periodic) in $x^{\prime}$, while it is constant in $y^{\prime}$ and $z^{\prime}$ due to the absence of spin-orbit coupling along these two directions. The quantity $k_{1}$ represents the half-length of the stripe wave vector, $2k_{1}\hat{\bm{e}}_{x}$. Its optimal value, along with the expansion coefficients $\tilde{\Psi}_{\bar{m}}$, is determined by minimizing the energy (3) for an order parameter of the form (14). This leads to the following expression for $k_{1}$ as a function of the $\tilde{\Psi}_{\bar{m}}$’s [58]:
$k_{1}=k_{R}\frac{\sum_{\text{{\hbox{\bar{m}}} odd}}\bar{m}\tilde{\Psi}_{\bar{m}}^{\dagger}\sigma_{z}\tilde{\Psi}_{\bar{m}}}{\sum_{\text{{\hbox{\bar{m}}} odd}}\bar{m}^{2}\tilde{\Psi}_{\bar{m}}^{\dagger}\tilde{\Psi}_{\bar{m}}}\,.$ (15)
Typically, the optimal $k_{1}$ is found to be approximately equal to the value $k_{1}^{\mathrm{SP}}$ obtained within the single-particle model, with corrections due to interparticle interactions [70, 58]. One can also explore configurations with non-optimal $k_{1}$, whose stability has been analyzed in both the supersolid [79] and non-supersolid [61] regimes.
In the previous work on the stationary stripe phase [76, 58] the energy minimization procedure was carried out under the constraint of fixed particle number (4), which leads to the normalization condition
¯$\sum_{\text{{\hbox{\bar{m}}} odd}}\tilde{\Psi}_{\bar{m}}^{\dagger}\tilde{\Psi}_{\bar{m}}=1$ (16)
@@
for the expansion coefficients. Regarding the momentum (5), the independence of the order parameter (14) from $y$ and $z$ immediately implies that the corresponding components $\langle p_{y}\rangle$ and $\langle p_{z}\rangle$ vanish. On the other hand, the result $\langle p_{x}\rangle=0$ in the stationary stripe phase can be understood through more subtle symmetry considerations. The time-independent Gross-Pitaevskii equation (8) with $\bm{v}=0$ is invariant under the action of the two operators $\sigma_{x}\mathcal{P}$ and $\sigma_{z}\mathcal{T}$, which combine spin rotation with parity ($\mathcal{P}$) and time reversal ($\mathcal{T}$), respectively. [^1] In a spin-$1/2$ system, the time-reversal operator is $\mathcal{T}=\mathrm{i}\sigma_{y}\mathcal{K}$ where $\mathcal{K}$ denotes complex conjugation. Thus, the order parameter resulting from the application of either of these two operators to Eq. (14) remains a solution of Eq. (8) and can differ from the original only by a global phase rotation and spatial translation. In terms of the expansion coefficients, this means that $\tilde{\Psi}_{-\bar{m}}^{*}=\sigma_{x}\tilde{\Psi}_{\bar{m}}$ up to a phase factor [76, 58]. Using this property, one can easily show that $\langle p_{x}\rangle$ vanishes when $\bm{v}=0$. Conversely, when the stripes are in motion, the action of $\sigma_{x}\mathcal{P}$ and $\sigma_{z}\mathcal{T}$ on Eq. (14) generates a stripe pattern moving with velocity $-\bm{v}$. In this case, the $x$ component of the momentum (5) is nonzero and is related to the expansion coefficients $\tilde{\Psi}_{\bar{m}}$ as follows:
¯$\sum_{\text{{\hbox{\bar{m}}} odd}}\bar{m}\tilde{\Psi}_{\bar{m}}^{\dagger}\tilde{\Psi}_{\bar{m}}=\frac{\langle p_{x}\rangle}{N\hbar k_{1}}\equiv s_{p}\,.$ (17)
Neglecting the typically small contribution from the $|\bar{m}|>1$ harmonics in Eq. (14), one can interpret $s_{p}$ as the imbalance between the occupations of the two single-particle energy minima. For a Bloch-wave-like order parameter of the form (14), the momentum per particle is constrained to the first Brillouin zone, meaning that $-\hbar k_{1}\leq\langle p_{x}\rangle/N\leq\hbar k_{1}$, which leads to the condition $-1\leq s_{p}\leq 1$. As we will see, $s_{p}$ in our spin-orbit-coupled BEC plays a role similar to that of the spin imbalance $\langle\sigma_{z}\rangle/N$ in the absence of spin-orbit coupling. In this analogy, $s_{p}=0$ corresponds to a balanced coherent mixture of the two degenerate minima of the single-particle spectrum, which defines the stripe phase at equilibrium. On the other hand, we will show that $s_{p}=1$ ($s_{p}=-1$) corresponds to a fully polarized state in momentum space, where all the atoms condense into the single-particle ground state with positive (negative) momentum. This is nothing else than the well-known plane-wave phase (see Sec. II). Interestingly, $s_{p}$ and $\langle\sigma_{z}\rangle/N$ coincide in the limit of zero Raman coupling, where the single-particle energy minima are characterized by well-defined spin (see Sec. IV.1). The explicit determination of the order parameter (14) for moving stripe patterns and the related observables will be discussed in the next section.
@@ IV Properties of moving striped patterns
After establishing the general framework, we now proceed with a detailed analysis of the moving stripe patterns. We begin by examining the exact solution for the case of zero Raman coupling, which can be mapped to an unbalanced BEC mixture (Sec. IV.1). Next, in Sec. IV.2, we describe the numerical and perturbative methods used to calculate the order parameter at finite Raman coupling. Finally, the results for several key observables are presented in Sec. IV.3.
@@
At zero Raman coupling ($\Omega_{R}=0$), the lowest-energy solution of Eq. (8) that satisfies the constraints (16) and (17) for a given $s_{p}$ has a simple analytical form:
$\Psi_{0}^{(0)}(\bm{r}^{\prime})=\sqrt{\bar{n}}\left[\begin{pmatrix}\tilde{\Psi}_{+1,\uparrow}^{(0)}\\ 0\end{pmatrix}\mathrm{e}^{\mathrm{i}k_{R}x^{\prime}}+\begin{pmatrix}0\\ \tilde{\Psi}_{-1,\downarrow}^{(0)}\end{pmatrix}\mathrm{e}^{-\mathrm{i}k_{R}x^{\prime}}\right]\,,$ (18)
@@
where $\tilde{\Psi}_{+1,\uparrow}^{(0)}=\sqrt{(1+s_{p})/2}\,\mathrm{e}^{\mathrm{i}(\theta+\Delta\theta/2)}$ and $\tilde{\Psi}_{-1,\downarrow}^{(0)}=\sqrt{(1-s_{p})/2}\,\mathrm{e}^{\mathrm{i}(\theta-\Delta\theta/2)}$. This solution has $\mu=G_{dd}$ and $\Delta\mu/2\equiv\hbar k_{R}v_{x}=G_{ss}s_{p}$, where $G_{dd}=g_{dd}\bar{n}$ and $G_{ss}=g_{ss}\bar{n}$. Equation (18) represents a special case of the stripe Ansatz (14), where only the $\bar{m}=\pm 1$ terms are nonzero, and the wave vector takes the value $k_{1}=k_{R}$, which results from energy minimization. In contrast, the global phase of the order parameter, $\theta$, and the relative phase between the two spatially oscillating terms, $\Delta\theta$, are not determined by energy minimization and can be chosen arbitrarily. The energy per particle for the configuration (18) in the laboratory frame is $E/N=(G_{dd}+G_{ss}s_{p}^{2})/2$. The relation $\langle\bm{p}\rangle=\hbar k_{R}\langle\sigma_{z}\rangle\hat{\bm{e}}_{x}$ holding at zero $\Omega_{R}$ implies that in this case, $s_{p}$ can be used to quantify both the spin polarization, $\langle\sigma_{z}\rangle=Ns_{p}$, and the momentum along $x$, $\langle p_{x}\rangle=N\hbar k_{R}s_{p}$. We emphasize that only the latter physical interpretation remains valid at finite Raman coupling [see Eq. (17)].
@@
Equation (18) describes a standard two-component BEC mixture with uniform density $\bar{n}$ and spin imbalance $\langle\sigma_{z}\rangle/N=s_{p}$. This becomes even more evident when considering the full order parameter in the laboratory frame, given by $\mathrm{e}^{-\mathrm{i}\mu t/\hbar}\Psi_{0}^{(0)}(\bm{r}-\bm{v}t)$ [see Eq. (7)], and performing the space-dependent spin rotation $\exp(-\mathrm{i}k_{R}\sigma_{z}x)$. This procedure transforms the order parameter into the canonical space-independent form [^2] Applying the unitary transformation $\exp(-\mathrm{i}k_{R}\sigma_{z}x)$ to the single-particle Hamiltonian (2) with $\Omega_{R}=\delta_{R}=0$ reduces it to the standard kinetic energy $\bm{p}^{2}/2m$. Consequently, Eq. (1) becomes the Gross-Pitaevskii equation for a two-component BEC without spin-orbit coupling, and its uniform solutions take the form in Eq. (19).
$\Psi_{\mathrm{mix}}(t)=\sqrt{\bar{n}}\begin{pmatrix}\tilde{\Psi}_{+1,\uparrow}^{(0)}\mathrm{e}^{-\mathrm{i}\mu_{\uparrow}t/\hbar}\\ \tilde{\Psi}_{-1,\downarrow}^{(0)}\mathrm{e}^{-\mathrm{i}\mu_{\downarrow}t/\hbar}\end{pmatrix}\,.$ (19)
@@
Notice that the combinations $\mu_{\uparrow,\downarrow}=\mu\pm\Delta\mu/2=G_{dd}\pm G_{ss}s_{p}$ correspond exactly to the chemical potentials of the two components in a binary BEC mixture with equal masses and symmetric intraspecies interactions [80, 81]. The global phase of the order parameter (19), $\theta-\mu t/\hbar$, evolves in time at a rate determined by the average chemical potential, $\mu=(\mu_{\uparrow}+\mu_{\downarrow})/2$. The freedom in choosing the value of $\theta$ at $t=0$ reflects the spontaneous breaking of global $\mathrm{U}(1)$ phase symmetry, a hallmark of Bose-Einstein condensation. The polarization vector of the mixture (19) is defined as $\langle\bm{\sigma}\rangle_{\mathrm{mix}}=\left(\langle\sigma_{x}\rangle_{\mathrm{mix}},\langle\sigma_{y}\rangle_{\mathrm{mix}},\langle\sigma_{z}\rangle_{\mathrm{mix}}\right)$, where
@@
$\langle\sigma_{x}\rangle_{\mathrm{mix}}$ ${}=N\sqrt{1-s_{p}^{2}}\cos(\Delta\mu\,t/\hbar-\Delta\theta)\,,$ (20a)
@@
$\langle\sigma_{y}\rangle_{\mathrm{mix}}$ ${}=N\sqrt{1-s_{p}^{2}}\sin(\Delta\mu\,t/\hbar-\Delta\theta)\,,$ (20b)
$\langle\sigma_{z}\rangle_{\mathrm{mix}}$ ${}=Ns_{p}\,,$ (20c)

### v2

@@ III.2 Traveling Bloch-wave Ansatz
where $\mu$ is the chemical potential and $\Psi_{0}$ is a function of the comoving coordinate xvtx-vt. Substituting the expression (17) into the Gross-Pitaevskii equation (1) and transforming to the new independent variables $x^{\prime}=x-vt$ and $t^{\prime}=t$ (i.e., the coordinates in the comoving frame, see Sec. V.1 below), we obtain the following time-independent equation for $\Psi_{0}$:
[…]
@@
where now the spatial derivatives are with respect to $x^{\prime}$. For $v=0$, Eq. (18) reduces to the standard time-independent Gross-Pitaevskii equation, whose solutions represent the stationary states of the condensate, including the ground state. The $v$-dependent term on the right hand side was previously included in Ref. [63] to construct and study plane-wave states with non-optimal momentum.
Having mapped our problem into a stationary one, we can now follow the same strategy as Refs. [78, 60] and look for a solution of Eq. (18) in the form of a Bloch wave, describing a configuration with periodic density modulations along $x$:
$\Psi_{0}(x^{\prime})=\mathrm{e}^{\mathrm{i}k_{c}x^{\prime}}\sum_{\text{$\bar{m}$ odd}}\tilde{\Psi}_{\bar{m}}\mathrm{e}^{\mathrm{i}\bar{m}k_{1}x^{\prime}}\,.$ (19)
@@
This order parameter is given by the product of a plane wave with wave vector $\bm{k}_{c}=k_{c}\hat{\bm{e}}_{x}$ (which can be identified with the mean condensation wave vector appearing within the two-harmonic approximation, see Sec. III.1) and of a $\pi/k_{1}$-antiperiodic (and therefore $2\pi/k_{1}$-periodic) function of $x^{\prime}$. This function is conveniently represented as an expansion in the basis of plane waves with wave vector equal to an half-integer multiple of the stripe wave vector $2\bm{k}_{1}=2k_{1}\hat{\bm{e}}_{x}$ [note that the sum in Eq. (19) runs over both positive and negative odd integers]. The corresponding two-component expansion coefficients are denoted by $\tilde{\Psi}_{\bar{m}}$. The two-harmonic Ansatz (11) can be recovered from Eqs. (17)–(19) by retaining only the $\bar{m}=\pm 1$ terms associated with the occupation of the two minima of the single-particle spectrum. In contrast, the additional $|\bar{m}|>1$ harmonics arise due to the nonlinearity of the Gross-Pitaevskii equation (18) that $\Psi_{0}$ satisfies. The mean chemical potential can be expressed in terms of $\mu$ and $v$ as $\mu_{d}=\mu+\hbar k_{c}v$, while for the chemical potential semi-difference one has $\mu_{s}=\hbar k_{1}v$, consistent with Eq. (16).
The average momentum density computed from the order parameter (19) is still given by Eq. (15), but the average particle and dressed spin polarization densities now read
$\bar{n}=\sum_{\text{$\bar{m}$ odd}}\tilde{\Psi}_{\bar{m}}^{\dagger}\tilde{\Psi}_{\bar{m}}$ (20)
and
$\bar{s}=\sum_{\text{$\bar{m}$ odd}}\bar{m}\tilde{\Psi}_{\bar{m}}^{\dagger}\tilde{\Psi}_{\bar{m}}\,$, (21)
respectively. Note that $\bar{s}$ enters the contribution of the antiperiodic part of the order parameter (19) to the average momentum density (15). We will take this contribution in the first Brillouin zone, i.e., between $\bar{n}\hbar k_{1}$ and $\bar{n}\hbar k_{1}$, which is always possible because $k_{c}$ is only defined up to a multiple of $2k_{1}$. Consequently, the population imbalance (14) ranges between 1-1 and $1$ even with the new definitions of $\bar{n}$ and $\bar{s}$.
@@
In the stationary stripe phase, the result $\bar{p}_{x}=0$ can be understood through symmetry considerations. The time-independent Gross-Pitaevskii equation (18) with $v=0$ is invariant under the action of the two operators $\sigma_{x}\mathcal{P}$ and $\sigma_{z}\mathcal{T}$, which combine spin rotation with parity ($\mathcal{P}$) and time reversal ($\mathcal{T}$), respectively. [^1] In a spin-$1/2$ system, the time-reversal operator is $\mathcal{T}=\mathrm{i}\sigma_{y}\mathcal{K}$ where $\mathcal{K}$ denotes complex conjugation. Thus, the order parameter resulting from the application of either of these two operators to Eq. (19) remains a solution of Eq. (18) and can differ from the original only by a global phase rotation and spatial translation. This can happen only if $k_{c}=0$ and the expansion coefficients fulfill, up to a phase factor, the equality $\tilde{\Psi}_{-\bar{m}}^{*}=\sigma_{x}\tilde{\Psi}_{\bar{m}}$ [78, 60]. Using these properties, one can easily show that $\bar{s}$, and thus $\bar{p}_{x}$, vanish when $v=0$. Conversely, when the stripes are in motion, the action of $\sigma_{x}\mathcal{P}$ and $\sigma_{z}\mathcal{T}$ on Eq. (19) generates a stripe pattern moving with velocity v-v along $x$. The corresponding momentum density (15) is therefore nonzero and changes sign when one of the two above symmetry operations is performed. This is achieved by simultaneously changing the sign of $k_{c}$ and $\bar{s}$.
The time-independent Gross-Pitaevskii equation (18) corresponds to the stationarity condition of the functional
$\begin{split}\bar{\varepsilon}_{\mathrm{GC}}&{}=\bar{\varepsilon}-\mu\bar{n}-v\bar{p}_{x}\\ &{}=\bar{\varepsilon}-\mu_{d}\bar{n}-\mu_{s}\bar{s}\,.\end{split}$ (22)
This functional can be regarded as the grand-canonical energy density of the system, with a constraint on the average particle density and another on the momentum or the dressed spin polarization density. Since $(\bar{n},\mu_{d})$ and $(\bar{s},\mu_{s})$ are pairs of conjugate thermodynamic variables, they fulfill the standard thermodynamic relations
$\mu_{d}={}$ ${}\frac{\partial\bar{\varepsilon}}{\partial\bar{n}}\,$, (23a)
$\mu_{s}={}$ ${}\frac{\partial\bar{\varepsilon}}{\partial\bar{s}}\,.$ (23b)
It is also useful to consider the derivatives of the average energy density with respect to $k_{c}$ and $k_{1}$. They are given by
$\bar{j}_{d,x}={}$ ${}\frac{\partial\bar{\varepsilon}}{\partial(\hbar k_{c})}=\frac{\hbar}{m}(k_{c}\bar{n}+k_{1}\bar{s}-k_{R}\bar{s}_{z})\,$, (24a)
$\bar{j}_{s,x}={}$ ${}\frac{\partial\bar{\varepsilon}}{\partial(\hbar k_{1})}=\frac{\hbar}{m}(k_{c}\bar{s}+k_{1}\bar{s}_{2}-k_{R}\bar{s}_{1z})\,$, (24b)
where $\bar{n}$ and $\bar{s}$ are as in Eq. (20) and (21), respectively, and
[…]
Specifically, Eqs. (24) are the derivatives of the kinetic energy density, which is the only term that depends explicitly on $k_{c}$ and $k_{1}$. Equation (24a) can be rewritten as $\bar{j}_{d,x}=\frac{1}{V}\int_{V}\mathrm{d}^{3}r\,j_{d,x}$, where $j_{d,x}$ is the $x$ component of the particle current density (8). Therefore, $\bar{j}_{d,x}$ is nothing else than the average particle current density along $x$. On the other hand, the physical interpretation of $\bar{j}_{s,x}$ is immediate in the limit of zero Raman coupling (see Sec. IV.1), where it reduces to the $x$ component of the average bare spin current density. We thus argue that, at finite $\Omega_{R}$, $\bar{j}_{s,x}$ plays the role of the average dressed spin current density in the $x$ direction. As discussed in Sec. III.3 below, in this work we will only consider configurations with optimal values of $k_{c}$ and $k_{1}$, and thus with zero $\bar{j}_{d,x}$ and $\bar{j}_{s,x}$. In particular, the vanishing of $\bar{j}_{d,x}$ entails that the momentum along $x$ is always proportional to the bare spin polarization, that is,
$\bar{p}_{x}=\hbar k_{R}\bar{s}_{z}\,.$ (26)
In addition, having $\bar{j}_{d,x}=0$ and $\bar{j}_{s,x}=0$ means that the stripe motion produces only oscillatory microscopic currents and no macroscopic mass transport. We mention that, for stationary stripes, configurations with non-optimal $k_{c}$ [66] or $k_{1}$ [82] have been explored, and found to be stable when the two wave vectors are sufficiently close to the optimal values.
III.3 Evaluation of the traveling stripe order parameter
Having presented the structure and the general properties of the order parameter of traveling stripe configurations, we now discuss how it can be explicitly determined. In the present work, we employ two complementary approaches. The first is a variational method, which begins inserting the Ansatz (19) into the total energy (3) and performing the spatial integration. One obtains an expression for the energy density $\bar{\varepsilon}=E/V$ as a function of $k_{c}$, $k_{1}$, and the two-component expansion coefficients $\tilde{\Psi}_{\bar{m}}$. Using this result in combination with Eqs. (20) and (21) one immediately determines the grand-canonical energy (22). We then minimize this energy with the Lagrange multipliers $\mu_{d}$ and $\mu_{s}$ chosen so that $\bar{n}$ and $\bar{s}$ take the desired values. As mentioned earlier, the stationarity condition of $\bar{\varepsilon}_{\mathrm{GC}}$ with respect to $k_{c}$ and $k_{1}$ implies the vanishing of the average particle and dressed spin current densities, see Eqs. (24). To perform the numerical minimization over the components of the $\tilde{\Psi}_{\bar{m}}$’s, the expansion (19) must be truncated to a finite number of harmonics. In our calculations, we retain terms with $9\leq\bar{m}\leq 9$, which is sufficient to achieve high accuracy in all physical quantities of interest. The validity of this variational method is confirmed by a direct comparison with the numerical solution of the Gross-Pitaevskii equation (18) in a periodic box. We find that the obtained wave function approaches the outcome of the variational calculation in the limit of large simulation boxes. [^2] We recall that, for a periodic box of length $L_{x}$ along $x$, the allowed values of $k_{c}$ and $k_{1}$ must be integer multiples of $\pi/L_{x}$. The spacing between these values decreases as $L_{x}$ increases and vanishes in the limit $L_{x}\to\infty$. In this limit, $k_{c}$ and $k_{1}$ become continuous variables, and their optimal values approach those obtained from the variational analysis. In particular, an examination of the momentum distribution reveals the presence of peaks associated with higher-order harmonics, thereby confirming the validity of including these terms in the Ansatz (19).
In Ref. [60], a perturbative approach was developed to analytically compute the equilibrium order parameter in the stripe phase, along with the associated observables. This method has proven highly accurate deeply in the double-minimum regime of the single-particle spectrum, which corresponds to low Raman coupling, i.e., $\hbar\Omega_{R}/4E_{R}\ll 1$. In the present work, we extend this perturbative approach to investigate the case of a moving striped condensate, thereby gaining deeper insight into the underlying physics. This method serves as an analytical alternative to the numerical minimization discussed earlier. Notably, the two approaches yield excellent agreement for small enough values of the Raman coupling (for the values of the interaction parameters we use in this work, we find that this agreement holds up to approximately $\hbar\Omega_{R}/E_{R}=1.5$). We begin by considering the power series expansion of the order parameter,
$\Psi_{0}=\Psi_{0}^{(0)}+\sum_{l=1}^{+\infty}\Psi_{0}^{(l)}\,$, (27)
along with the expansions for the mean chemical potential,
$\mu_{d}=\mu_{d}^{(0)}+\sum_{l=1}^{+\infty}\mu_{d}^{(l)}\,$, (28)
and the chemical potential semi-difference,
$\mu_{s}=\mu_{s}^{(0)}+\sum_{l=1}^{+\infty}\mu_{s}^{(l)}\,.$ (29)
Here and in the following, the superscript “(l)(l)” denotes the contribution of order $l$ in the small expansion parameter $\hbar\Omega_{R}/4E_{R}$. The zero order terms in each series coincide with the results obtained in the absence of Raman coupling (see Sec. IV.1 and Appendix A). Substituting these expansions into the stationary Gross-Pitaevskii equation (18) yields a set of recurrence relations, which allow for the computation of each order-$l$ correction in terms of lower-order results up to l1l-1. A detailed derivation of this procedure, together with the expressions for the order parameter up to the second order in $\hbar\Omega_{R}/4E_{R}$, is given in Appendix A and Appendix B.
@@ IV Properties of moving striped patterns
After establishing the general framework, we now proceed with a detailed analysis of the moving stripe patterns. We begin by examining the exact solution for the case of zero Raman coupling, which can be mapped to an unbalanced BEC mixture (Sec. IV.1). Then, we present the results for several key observables at finite Raman coupling (Sec. IV.2).
@@
At zero Raman coupling ($\Omega_{R}=0$), the lowest-energy solution of Eq. (18) that satisfies the constraints (20) and (21) has a simple analytical form:
$\Psi_{0}^{(0)}(x^{\prime})=\begin{pmatrix}\tilde{\Psi}_{+1,\uparrow}^{(0)}\\ 0\end{pmatrix}\mathrm{e}^{\mathrm{i}k_{R}x^{\prime}}+\begin{pmatrix}0\\ \tilde{\Psi}_{-1,\downarrow}^{(0)}\end{pmatrix}\mathrm{e}^{-\mathrm{i}k_{R}x^{\prime}}\,$, (30)
@@
where $\tilde{\Psi}_{+1,\uparrow}^{(0)}=\sqrt{\bar{n}(1+s_{p})/2}\,\mathrm{e}^{\mathrm{i}(\theta+\Delta\theta/2)}$ and $\tilde{\Psi}_{-1,\downarrow}^{(0)}=\sqrt{\bar{n}(1-s_{p})/2}\,\mathrm{e}^{\mathrm{i}(\theta-\Delta\theta/2)}$. This solution has $\mu_{d}^{(0)}=G_{dd}$ and $\mu_{s}^{(0)}=G_{ss}s_{p}$, where $G_{dd}=g_{dd}\bar{n}$ and $G_{ss}=g_{ss}\bar{n}$. Equation (30) represents a special case of the stripe Ansatz (19), where only the $\bar{m}=\pm 1$ terms are nonzero. The optimal values of the wave vectors, $k_{c}^{(0)}=0$ and $k_{1}^{(0)}=k_{R}$, result from energy minimization. In contrast, the global phase of the order parameter, $\theta$, and the relative phase between the two spatially oscillating terms, $\Delta\theta$, are not determined by energy minimization and can be chosen arbitrarily. The energy density for the configuration (30) is $\bar{\varepsilon}^{(0)}=\bar{n}(G_{dd}+G_{ss}s_{p}^{2})/2$.
@@
Equation (30) describes a standard two-component BEC mixture with uniform particle density $\bar{n}$ and bare spin imbalance $\bar{s}_{z}^{(0)}/\bar{n}=s_{p}$ (note that dressed spins reduce to their bare counterparts when $\Omega_{R}=0$). This becomes even more evident when considering the full order parameter, given by $\mathrm{e}^{-\mathrm{i}\mu t/\hbar}\Psi_{0}^{(0)}(x-vt)$ [see Eq. (17)], and performing the space-dependent spin rotation $\exp(-\mathrm{i}k_{R}\sigma_{z}x)$. This procedure transforms the order parameter into the canonical space-independent form [^3] Applying the unitary transformation $\exp(-\mathrm{i}k_{R}\sigma_{z}x)$ to the single-particle Hamiltonian (2) with $\Omega_{R}=\delta_{R}=0$ reduces it to the standard kinetic energy $\bm{p}^{2}/2m$. Consequently, Eq. (1) becomes the Gross-Pitaevskii equation for a two-component BEC without spin-orbit coupling, and its uniform solutions take the form in Eq. (31).
$\Psi_{\mathrm{mix}}(t)=\begin{pmatrix}\tilde{\Psi}_{+1,\uparrow}^{(0)}\mathrm{e}^{-\mathrm{i}\mu_{\uparrow}t/\hbar}\\ \tilde{\Psi}_{-1,\downarrow}^{(0)}\mathrm{e}^{-\mathrm{i}\mu_{\downarrow}t/\hbar}\end{pmatrix}\,$. (31)
@@
Notice that the combinations $\mu_{\uparrow,\downarrow}=\mu_{d}^{(0)}\pm\mu_{s}^{(0)}=G_{dd}\pm G_{ss}s_{p}$ correspond exactly to the chemical potentials of the two components in a binary BEC mixture with equal masses and symmetric intraspecies interactions [83, 84]. The global phase of the order parameter (31), $\theta-\mu_{d}^{(0)}t/\hbar$, evolves in time at a rate determined by the mean chemical potential, $\mu_{d}^{(0)}=(\mu_{\uparrow}+\mu_{\downarrow})/2$. The freedom in choosing the value of $\theta$ at $t=0$ reflects the spontaneous breaking of global $\mathrm{U}(1)$ phase symmetry, a hallmark of Bose-Einstein condensation. The polarization density vector of the mixture (31) is defined as $\langle\bm{\sigma}\rangle_{\mathrm{mix}}/V=\left(\langle\sigma_{x}\rangle_{\mathrm{mix}},\langle\sigma_{y}\rangle_{\mathrm{mix}},\langle\sigma_{z}\rangle_{\mathrm{mix}}\right)/V$, where
@@
$\frac{\langle\sigma_{x}\rangle_{\mathrm{mix}}}{V}$ ${}=\bar{n}\sqrt{1-s_{p}^{2}}\cos\left[2\mu_{s}^{(0)}t/\hbar-\Delta\theta\right]\,$, (32a)
@@
$\frac{\langle\sigma_{y}\rangle_{\mathrm{mix}}}{V}$ ${}=\bar{n}\sqrt{1-s_{p}^{2}}\sin\left[2\mu_{s}^{(0)}t/\hbar-\Delta\theta\right]\,$, (32b)
$\frac{\langle\sigma_{z}\rangle_{\mathrm{mix}}}{V}$ ${}=\bar{n}s_{p}\,$, (32c)